sir generally if an equation has two logarathimic variables with different bases how to solve it


example


log4 (x-10= log2 (x-3)


where 4 and 2 are bases

Grade:12

1 Answers

vikas askiitian expert
13 years ago

loga b = logb/loga

 

log4(x-10) = log2(x-3)

log(x-4)/log4 = log(x-3)/log2

{ log4 = log22 = 2log2 }

log(x-4)/2log2 = log(x-3)/log2

log(x-4) = 2log(x-3) = log(x-3)2

x-4 = (x-3) 

x2 -7x + 13 = 0

now solve this & get x

 

approve if u like my ans